OpenStudy (anonymous):

Solve 3x2 − 3x − 6 = 0

OpenStudy (anonymous):

I did this 3x2 − 3x − 6 = 0. + 6 = 6 3x2 - 3x = 6 / +3 /3 3x2 - x = 2 \3 - x = 3 x2 - x = 2/3

OpenStudy (anonymous):

But the answer must be There : {1, 2} {-2, 1} {-1, 2} {-2, -1}

OpenStudy (zzr0ck3r):

3x2 − 3x − 6 = 0 a=3, b=-3, c=-6 a*c = -18 we want two numbers that multiply to -16 but add to -3 -6 and 3 are these numbers expand the middle term 3x^2+3x-6x-6=0 3x(x+1)-6(x+1)=0 (3x-6)(x+1)=0 so this is zero when 3x-6=0 and x+1=0 so 3x-6=0 3x=6 x=2 and x+1=0 x=-1 so {-1,2} is the solution set.

OpenStudy (zzr0ck3r):

you went wrong here 3x2 - 3x = 6 / +3 /3